OpenStudy (rainbow_dashie):

At noon, the temperature of a container of ether was -109C. At 1 p.m., the temperature of the container of ether was -34C. What was the change in temperature from noon to 1 p.m.?

OpenStudy (rainbow_dashie):

@kevindragonlord

OpenStudy (anonymous):

There are many ways to write the abbreviation for "post meridiem." It can be in capital letters, small capitals, or lowercase, with or without the periods

OpenStudy (anonymous):

it is most likely 109c

OpenStudy (dumbcow):

temp increased from -109 to -34 the change is simply the difference -----> (-34 - (-109)) = 109 -34 = 75